OpenStudy (anonymous):

can someone explain to me how to do this?? The Simanski family is purchasing a new sofa. They find one they like that is on sale for 15% off. If the sales tax rate in their state is 6% , what will be the final cost on the sofa that was originally priced at 1,200 ?

OpenStudy (anonymous):

\[1200\times .85\times (1.06)\]

OpenStudy (anonymous):

75/100*106/100*1200=954

OpenStudy (anonymous):

@ashwinjohn3 15% off, not 25% off

OpenStudy (anonymous):

Ohh thanks 85/100*106/100*1200=1 081.2